OpenStudy (anonymous):

How to convert 2x^2−44x+246 to vertex form?

OpenStudy (anonymous):

@triciaal Here it is :o but I have to get to bed right now, so it would be great if you could tell me the steps and ill look at them when I get back on, tomorrow ;w; thanks again! XD

OpenStudy (triciaal):

it's basically completing the square then factor of 2 start with ax^2 + bx + c = 0 first divide by a to work with coefficient of x^2 = 1 apply (a -b)^2 = a^2 -2ab + b^2

OpenStudy (triciaal):

do you have enough to sleep on?

OpenStudy (triciaal):